OpenStudy (anonymous):

Given the line; 2x-3y=5, find the following. a. A vector that is parallel to the line. b. A vector that is perpendicular to the line.

OpenStudy (cuanchi):

if you convert the line in the form of f(x)=ax+b where a is the slope of the line, and b is the value of y for x=0 you get from 2x-3y=5 => y=(2/3)x-(5/3) any other function with the same value of a and different value of b will be parallel to 2x-3y=5 then if you change the sign of the a you will have a perpendicular line to 2x-3y=5